# Extracting the Luttinger parameter from a single wave function

### v1

@@ Abstract
with $K=1/2$ that provides a perfect lattice realization of the SU(2)${}_{1}$ Wess-Zumino-Witten model [58, 59]. Both models have a U(1) symmetry associated with the conservation of the $z$-component of the total spin.
The success of Eq. (4) suggests that
$|\mathcal{C}_{\mathrm{latt}}\rangle=\prod_{j=1}^{N/2}\left(|\uparrow\rangle_{j}|\uparrow\rangle_{j+N/2}+|\downarrow\rangle_{j}|\downarrow\rangle_{j+N/2}\right)$ (13)
is likely a crosscap state for spin-1/2 chains with even $N$. However, there is one subtlety in this definition: it corresponds to $|\mathcal{C}_{+}\rangle$ when $\mathrm{mod}(N,4)=0$ but does not give the expected result when $\mathrm{mod}(N,4)=2$. The reason for this can be traced to the constraint $(S_{j}^{z}-S_{j+N/2}^{z})|\mathcal{C}_{\mathrm{latt}}\rangle=0$. Using bosonization, $S^{z}_{j}$ operator reads $S^{z}_{j}\sim\partial_{x}\hat{\theta}(x)+(-1)^{j}\alpha\cos(\hat{\theta}(x)/R^{\prime})$ for XXZ and HS chains, where $\alpha$ is a model-dependent constant. For $\mathrm{mod}(N,4)=0$, the lattice constraint $(S_{j}^{z}-S_{j+N/2}^{z})|\mathcal{C}_{\mathrm{latt}}\rangle=0$ is hence consistent with the constraints of $|\mathcal{C}_{+}\rangle$ in Eq. (6).
We now proceed to demonstrate the validity of Eq. (13). For the XY model (i.e., $\Delta=0$ in the XXZ model) and the HS model, explicit wave functions of many eigenstates are known [60]. This allows us to derive analytical results to support our theory. In both models, the eigenstates that correspond to the Virasoro primary states $|n,0\rangle$ have a Jastrow form in the hardcore boson basis:
$\Psi^{\lambda}(x_{1},\ldots,x_{M})\propto(-1)^{\sum_{l}x_{l}}\prod_{j<k}\left|\sin\frac{\pi(x_{j}-x_{k})}{N}\right|^{\lambda}$ (14)
with $1\leq x_{1}<\cdots<x_{M}\leq N$ denoting the positions of spin-$\uparrow$ sites (with hardcore bosons occupied), and $\lambda=1$ ($\lambda=2$) for the XY (HS) chain. The hardcore boson number $M$ is related to the U(1) charge $n$ via $M=N/2+n$. As $|\mathcal{C}_{\mathrm{latt}}\rangle$ contains only states with even number of bosons, $\langle\mathcal{C}_{\mathrm{latt}}|\Psi^{\lambda}(M)\rangle$ vanishes identically for odd $M$. This implies $\langle\mathcal{C}_{\mathrm{latt}}|\Psi^{\lambda}(M=N/2+2n)\rangle=0$ when $\mathrm{mod}(N,4)=2$. As $|\Psi^{\lambda}(M=N/2+2n)\rangle$ corresponds to the Virasoro primary state $|2n,0\rangle$, this indicates $|\mathcal{C}_{\mathrm{latt}}\rangle$ does not realize $|\mathcal{C}_{\pm}\rangle$ in this case [see Eq. (10)]. For $\mathrm{mod}(N,4)=0$, we have $\langle\mathcal{C}_{\mathrm{latt}}|\Psi^{\lambda}(M)\rangle=0$ for odd $M=N/2+2n+1$ and
@@
for even $M=N/2+2n$ [56]. This is in perfect agreement with the field theory prediction in Eq. (10), giving a strong indication that $|\mathcal{C}_{\mathrm{latt}}\rangle$ corresponds to $|\mathcal{C}_{+}\rangle$ for $\mathrm{mod}(N,4)=0$.
@@
For the XXZ model with $\Delta{\neq}0$, the ground state can be computed numerically using the density matrix renormalization group (DMRG) [61]. When the ground state $|\Psi\rangle$ is expressed as a matrix product state (MPS) [62, 63], its crosscap overlap $\langle\mathcal{C}_{\rm latt}|\Psi\rangle$ is given by the tensor contraction shown in Fig. 1(b). Numerical results at many different $N$’s and $\Delta$’s are displayed in Fig. 2. The MPS bond dimension is $2000$ for $N=100$ and $6000$ for $N=400$. The accuracy is very good in a wide range of $\Delta$ for a moderate $N=100$. The relative error of $K$ is of the order $10^{-4}$ or smaller when $\Delta\in[-0.95,0.8]$. As $\Delta$ approaches $1$, clear deviation from exact values is observed (not visible in Fig. 2(a) due to its scale) and the relative error reaches $7.20\%$ at $\Delta=1$.
We attribute the obvious deviation at $\Delta=1$ to the marginal terms in the low-energy theory [64], which usually cause strong finite-size effects. The deviation gets smaller as $N$ increases, but the convergence towards the exact value is very slow [see Fig. 2(b)]. To validate our conjecture, we add a next-nearest-neighbor coupling to the XXZ model with $\Delta=1$:
$H_{\rm next}=\sum_{j=1}^{N}(\mathbf{S}_{j}\cdot\mathbf{S}_{j+1}+J_{2}\mathbf{S}_{j}\cdot\mathbf{S}_{j+2})\,.$ (16)
Its low-energy physics contains no marginal term at $J_{2}{\approx}0.24$ [65]. The results for $N=100$ are presented in Fig. 2(c). As $J_{2}$ increases toward $0.24$, the relative error of $K$ decreases and eventually becomes $0.11\%$. This confirms that the marginal terms are responsible for the deviation. Further analysis based on conformal perturbation theory suggests that the Luttinger parameter should be fitted as $K=a[\ln(N)]^{-2}+b$ [57]. The data points with $N\in[100,400]$ in Fig. 2(b) thus yield $K=0.523$ in the thermodynamic limit.
@@
$H_{q{\rm SC}}=-\sum_{j=1}^{N}(\sigma^{{\dagger}}_{j}\sigma_{j+1}+\sigma_{j}\sigma^{{\dagger}}_{j+1})-h\sum_{j=1}^{N}(\tau_{j}+\tau_{j}^{{\dagger}})\,,$ (17)
where the $\mathbb{Z}_{q}$ spin operators are defined by $\sigma|\alpha\rangle=|{\rm mod}(\alpha-1+q,q)\rangle$ and $\tau|\alpha\rangle=e^{i2\pi\alpha/q}|\alpha\rangle$ with $\alpha=1,\ldots,q$. This model is self-dual (SD) at $h=1$. For $q{\geq}5$ and intermediate $h$ (including the SD point), this model hosts a Luttinger liquid that is separated by two Kosterlitz–Thouless (KT) transitions at $h_{\rm KT1}$ and $h_{\rm KT2}$ from gapped phases at small and large $h$ [72, 73]. The transition points are related to each other by duality via $h_{\rm KT1}h_{\rm KT2}=1$. There is no analytical expression for the Luttinger parameter, but its value at the SD point and the KT transition points was predicted to be [68, 34]
$K_{\rm KT1}=\frac{q^{2}}{8},\quad K_{\rm SD}=\frac{q}{2},\quad K_{\rm KT2}=2\,.$ (18)
For the $q$-state clock model, we shall employ the lattice crosscap state $|\mathcal{C}_{\mathrm{latt}}\rangle=\prod_{j=1}^{N/2}\left(\sum_{\alpha=1}^{q}|\alpha\rangle_{j}|\alpha\rangle_{j+N/2}\right)$. Numerical results for the $q=5$ model with $N=50,60,\cdots,90$ and various $h$ are presented in Fig. 2(d). The MPS bond dimension is $1500$ for $N=50$ and $2500$ for $N=90$. At the SD point, finite-size systems are described by CFT with irrelevant perturbations, and the Luttinger parameter has a relative error $0.79\%$ at $N=90$. In the Luttinger liquid regime, the crosscap overlap only exhibits weak dependence on the system size, so the data points for different $N$ almost coincide. For the small and large $h$ phases, the overlap strongly depends on $N$. This picture is supported by Fig. 2(d), but the transition points cannot be located accurately due to the marginal terms at the KT transition points.
Summary and outlook — To summarize, we have demonstrated that the Luttinger parameter can be extracted from the overlaps between individual eigenstates and the crosscap states. The advantages of our method are corroborated by analytical and numerical results. It is also noteworthy that our method does not necessarily require a Hamiltonian, so it may be applied to check if a trial wave function indeed describes the physics of TLL or other CFTs (see, e.g., Refs. [74, 75]). From the theoretical perspective, our results open up an exciting new avenue in extracting conformal data and many interesting directions are yet to be explored. It is clear that the accuracy of our results is strongly affected by marginal terms, so one would like to improve the analytical formula to take their impact into account [57]. The crosscap state used here is constructed from a certain type of Bell states. For integrable spin chains, other types of crosscap states have been studied [47]. Do they have clear counterparts in CFTs? An extension of our method to include multi-component Luttinger liquids, such as the fermionic Hubbard model and SU(3) spin chains, is certainly very desirable. We hope that interesting results on these topics will be reported in the future.
@@ Appendix A Appendix A: Tomonaga-Luttinger liquid and compactified boson conformal field theory
The bosonized version of the (single-component) Tomonaga-Luttinger liquid is formulated as a free massless compactified boson field theory in (1+1)-dimensional spacetime. This section is devoted to a brief review of the operator formalism of the compactified boson field theory, and we mostly follow the notation used in Refs. [55, 76]. We assume that the spatial direction is compact with length $L$, and the system admits periodic boundary condition along this direction. Therefore, the spacetime manifold (or “world sheet”), parametrized by coordinates $(x~{}\mathrm{mod}~{}L,t)$, is effectively a cylinder of circumference $L$. The Hamiltonian reads
@@ Appendix B Appendix B: Determining crosscap states from constraints
up to overall factors $\kappa_{\pm}$. Without loss of generality, we choose $\kappa_{\pm}\in\mathbb{R}$ and $\kappa_{\pm}>0$ below.
@@ C.1 C1: XY chain
C.2 C2: Haldane-Shastry chain
@@
$|\tilde{\mathcal{N}}_{N,M,\lambda=2}|^{-2}=\frac{1}{M!}\cdot N^{M}\cdot\frac{(2M)!}{2^{M}}=\left(\frac{N}{2}\right)^{M}\frac{(2M)!}{M!}\,,$ (A63)

### v2

@@ Abstract
with $K=1/2$ [64, 65]. Both models have a U(1) symmetry associated with the conservation of the $z$-component of the total spin.
It is natural to speculate that
$|\mathcal{C}_{\mathrm{latt}}\rangle=\prod^{N/2}_{j=1}\left(|\uparrow\rangle_{j}|\uparrow\rangle_{j+N/2}+|\downarrow\rangle_{j}|\downarrow\rangle_{j+N/2}\right)$ (14)
is a crosscap state when $N$ is even [52, 53, 54, 55, 56], but there is one subtlety: it corresponds to $|\mathcal{C}_{+}\rangle$ only if $\mathrm{mod}(N,4)=0$ but does not for the cases with $\mathrm{mod}(N,4)=2$. This property can be traced back to the constraint $(S^{z}_{j}-S^{z}_{j+N/2})|\mathcal{C}_{\mathrm{latt}}\rangle=0$. The $S^{z}_{j}$ operator in bosonization reads $S^{z}_{j}\sim\partial_{x}\hat{\theta}(x)+(-1)^{j}\alpha\cos[\hat{\theta}(x)/R^{\prime}]$ for XXZ and HS chains ($\alpha$ is a model-dependent constant). For $\mathrm{mod}(N,4)=0$, the lattice constraint $(S^{z}_{j}-S^{z}_{j+N/2})|\mathcal{C}_{\mathrm{latt}}\rangle=0$ is hence consistent with Eq. (7).
The validity of Eq. (14) in the XY model (i.e., $\Delta=0$ in the XXZ model) and the HS model can be substantiated by analytical calculations. In the hardcore boson basis, the Virasoro primary states $|n,0\rangle$ have the wave functions $\Psi^{\lambda}(x_{1},\ldots,x_{M})\propto(-1)^{\sum_{l}x_{l}}\prod_{j<k}\left|\sin\frac{\pi(x_{j}-x_{k})}{N}\right|^{\lambda}$, where $1\leq x_{1}<\cdots<x_{M}\leq N$ denote the positions of spin-$\uparrow$ sites (occupied by hardcore bosons), and $\lambda=1$ ($\lambda=2$) for the XY (HS) chain [66]. The number of bosons $M$ is related to the U(1) charge $n$ via $M=N/2+n$. It is important to note that the cases with $\mathrm{mod}(N,4)=2$ and $\mathrm{mod}(N,4)=0$ are different. Because $|\mathcal{C}_{\mathrm{latt}}\rangle$ only contains states with even $M$, $\langle\mathcal{C}_{\mathrm{latt}}|\Psi^{\lambda}(M)\rangle$ vanishes identically for odd $M$. We also have $\langle\mathcal{C}_{\mathrm{latt}}|\Psi^{\lambda}(M=N/2+2n)\rangle=0$ when $\mathrm{mod}(N,4)=2$, so $|\mathcal{C}_{\mathrm{latt}}\rangle$ does not correspond to $|\mathcal{C}_{\pm}\rangle$ in these cases. In contrast, it can be proven that
@@
if $\mathrm{mod}(N,4)=0$ and $M=N/2+2n$ [62]. This is a perfect agreement with the field theory prediction Eq. (11), giving a strong indication that $|\mathcal{C}_{\mathrm{latt}}\rangle$ corresponds to $|\mathcal{C}_{+}\rangle$ for $\mathrm{mod}(N,4)=0$.
@@
For the XXZ model with $\Delta{\neq}0$, the ground state can be computed numerically using the density matrix renormalization group (DMRG) [67]. When the ground state $|\Psi\rangle$ is expressed as a matrix product state (MPS) [68, 69], its crosscap overlap $\langle\mathcal{C}_{\rm latt}|\Psi\rangle$ would be the tensor contraction shown in Fig. 1(b). Numerical results at many different $N$’s and $\Delta$’s are displayed in Fig. 2. The MPS bond dimension is $2000$ for $N=100$ and $6000$ for $N=400$. The accuracy is very good in a wide range of $\Delta$ for a moderate $N=100$. The relative error of $K$ is of the order $10^{-4}$ or smaller when $\Delta\in[-0.95,0.8]$. As $\Delta$ approaches $1$, clear deviation from exact values is observed [not visible in Fig. 2(a) due to its scale] and the relative error reaches $7.20\%$ at $\Delta=1$. This deviation is tentatively attributed to the marginal terms in the low-energy theory at $\Delta=1$ [70], which usually cause strong finite-size effects. It does get smaller as $N$ increases but the convergence is very slow [see Fig. 2(b)]. To validate this conjecture, we study the model $H_{\rm next}=\sum^{N}_{j=1}(\mathbf{S}_{j}\cdot\mathbf{S}_{j+1}+J_{2}\mathbf{S}_{j}\cdot\mathbf{S}_{j+2})$ with a next-nearest-neighbor Heisenberg interaction, for which the marginal term disappears when $J_{2}{\approx}0.24$ [71]. The results for this model with $N=100$ are presented in Fig. 2(c). As $J_{2}$ increases toward $0.24$, the relative error of $K$ decreases and eventually becomes $0.11\%$. Further analysis based on conformal perturbation theory reveals that finite-size results for the XXZ model with $\Delta=1$ should be fitted as $K=a[\ln(N)]^{-1}+b$ [62, 63]. This leads to $b=0.510$ using the data points with $N\in[100,400]$ in Fig. 2(b).
@@
$H_{q{\rm SC}}=-\sum^{N}_{j=1}(\sigma^{{\dagger}}_{j}\sigma_{j+1}+\sigma_{j}\sigma^{{\dagger}}_{j+1})-h\sum^{N}_{j=1}(\tau_{j}+\tau_{j}^{{\dagger}})\,,$ (16)
where the $\mathbb{Z}_{q}$ spin operators are defined as $\sigma|\alpha\rangle=|{\rm mod}(\alpha-1+q,q)\rangle$ and $\tau|\alpha\rangle=e^{i2\pi\alpha/q}|\alpha\rangle$ with $\alpha=1,\ldots,q$. This model is self-dual (SD) at $h=1$. For $q{\geq}5$ and intermediate $h$ (including the SD point), it hosts a Luttinger liquid that is separated by two Kosterlitz–Thouless (KT) transitions at $h_{\rm KT1}$ and $h_{\rm KT2}$ from gapped phases [78, 79]. The transition points are related to each other by duality as $h_{\rm KT1}h_{\rm KT2}=1$. There is no analytical expression for the Luttinger parameter, but its value at the SD point and the KT transition points was predicted to be $K_{\rm KT1}=q^{2}/8,K_{\rm SD}=q/2,K_{\rm KT2}=2$ [74, 40]. As in previous models, the Ansatz
$|\mathcal{C}_{\mathrm{latt}}\rangle=\prod_{j=1}^{N/2}\sum_{\alpha=1}^{q}|\alpha\rangle_{j}|\alpha\rangle_{j+N/2}$ (17)
is proposed as a crosscap state. Numerical results for the $q=5$ case with $N=50,\cdots,90$ and various $h$ are presented in Fig. 2(d). The MPS bond dimension is $1500$ for $N=50$ and $2500$ for $N=90$. At the SD point, finite-size systems are described by CFT with irrelevant perturbations, and the Luttinger parameter has a relative error $0.79\%$ at $N=90$. In the Luttinger liquid regime, the crosscap overlap only exhibits weak dependence on the system size, so the data points for different $N$ almost coincide. For the small and large $h$ phases, the overlap strongly depends on $N$.
Summary and outlook — To summarize, we have unveiled that the Luttinger parameter can be extracted from the overlaps between individual eigenstates and the crosscap states. The advantages of our method are corroborated by analytical and numerical results. This method may be applied to check if a trial wave function indeed describes the physics of TLL or other CFTs (see, e.g., Refs. [80, 81]) in the absence of a Hamiltonian. From the theoretical perspective, our results open up an exciting new avenue and many interesting directions are yet to be explored. The crosscap state used here is constructed from a certain type of Bell states. For integrable spin chains, other types of crosscap states have been studied [53]. Do they have clear counterparts in CFTs? An extension of our method to multicomponent Luttinger liquids, such as the fermionic Hubbard model and SU(3) spin chains, is certainly very desirable. Finally, experimental protocols for measuring the crosscap overlap should be pursued. We hope that interesting results on these topics will be reported in the future.
@@ Appendix A Appendix A: Tomonaga-Luttinger liquid and compactified boson conformal field theory
The bosonized version of the (single-component) Tomonaga-Luttinger liquid is formulated as a free massless compactified boson field theory in (1+1)-dimensional spacetime. This section is devoted to a brief review of the operator formalism of the compactified boson field theory, and we mostly follow the notation used in Refs. [61, 82]. We assume that the spatial direction is compact with length $L$, and the system admits periodic boundary condition along this direction. Therefore, the spacetime manifold (or “world sheet”), parametrized by coordinates $(x,t)$, is effectively a cylinder of circumference $L$. The Hamiltonian reads
@@ Appendix B Appendix B: Determining crosscap states from constraints
up to overall factors $\kappa_{\pm}$. Without loss of generality, we choose $\kappa_{\pm}>0$ below.
@@ C.1 XY chain
C.2 Haldane-Shastry chain
@@
$|\tilde{\mathcal{N}}_{N,M,\lambda=2}|^{-2}=\frac{1}{M!}\cdot N^{M}\cdot\frac{(2M)!}{2^{M}}=\left(\frac{N}{2}\right)^{M}\frac{(2M)!}{M!}\,.$ (A53)
@@
Appendix D Appendix D: Logarithmic correction to the crosscap overlap at the Heisenberg point
In this section, we provide a field theoretical analysis for understanding the finite-size scaling behavior of the overlap between the (lattice) crosscap state $|\mathcal{C}_{\mathrm{latt}}\rangle$ and the ground state of the spin-1/2 Heisenberg chain. In particular, we focus on the influence of the marginally irrelevant term.
The effective theory of the spin-1/2 Heisenberg chain is the $\mathrm{SU}(2)_{1}$ Wess-Zumino-Witten (WZW) CFT, with an SU(2) symmetric marginal perturbation:
$H=H_{0}+H_{1}$ (A56)
with
$H_{0}$ $=\frac{v}{2\pi}\int_{0}^{L}\mathrm{d}x\left[\frac{1}{3}:\mathbf{J}\cdot\mathbf{J}:(x)+\frac{1}{3}:\bar{\mathbf{J}}\cdot\bar{\mathbf{J}}:(x)\right]-\frac{2\pi v}{L}\frac{c}{12}\,,$
$H_{1}$ $=-vg\int_{0}^{L}\mathrm{d}x\;\frac{2}{\sqrt{3}}:\mathbf{J}\cdot\bar{\mathbf{J}}:(x)\,,$ (A57)
where $\mathbf{J}=(J^{0},J^{1},J^{2})$ and $\bar{\mathbf{J}}=(\bar{J}^{0},\bar{J}^{1},\bar{J}^{2})$ are the $\mathrm{SU}(2)_{1}$ Kac-Moody chiral and anti-chiral currents, respectively. Here, $v$ is the velocity, $c=1$ is the central charge, and $g$ is the coupling of the marginal term.
In the free-field representation, the $\mathrm{SU}(2)_{1}$ Kac-Moody currents are expressed in terms of the compactified boson field $\hat{\varphi}=\phi+\bar{\phi}$ with compactification radius $R=\sqrt{2}$:
$J^{0}(w)$ $=\frac{i}{\sqrt{2}}\partial\phi(w)\,,\quad J^{\pm}(w)=J^{1}(w)\pm iJ^{2}(w)=:e^{\pm i\sqrt{2}\phi(w)}:\,,$
$\bar{J}^{0}(\bar{w})$ $=\frac{i}{\sqrt{2}}\bar{\partial}\bar{\phi}(\bar{w})\,,\quad\bar{J}^{\pm}(\bar{w})=\bar{J}^{1}(\bar{w})\pm i\bar{J}^{2}(\bar{w})=:e^{\pm i\sqrt{2}\bar{\phi}(\bar{w})}:\,,$ (A58)
where $w=e^{\frac{2\pi}{L}(v\tau+ix)}$ is the complex coordinate on the plane. The unperturbed Hamiltonian $H_{0}$, corresponding to the $\mathrm{SU}(2)_{1}$ WZW CFT, represents a Tomonaga-Luttinger liquid with Luttinger parameter $K=R^{2}/4=1/2$. The effective coupling $g>0$ in $H_{1}$ can be understood from the bosonization treatment of the spin-1/2 Heisenberg chain, $S_{j}^{z}S_{j+1}^{z}\sim(\partial_{x}\hat{\vartheta}(x))^{2}:-\alpha^{2}\left(\frac{2\pi}{L}\right)^{2}:\cos[\sqrt{2}\hat{\vartheta}(x)]:$, with $\alpha$ being a non-universal constant.
Denoting the ground state of effective Hamiltonian [Eq. (A56)] as $|\Psi_{0}(g)\rangle$, its overlap with the conformal crosscap state $|\mathcal{C}_{+}\rangle$ [Eq. (A25) with $\kappa_{+}=2^{1/4}$ for the $\mathrm{SU}(2)_{1}$ WZW CFT], which has been identified as the continuum counterpart of the lattice crosscap state $|\mathcal{C}_{\mathrm{latt}}\rangle$, should be a universal scaling function of $g$. This overlap, denoted as $|\langle\mathcal{C}_{+}|\Psi_{0}(g)\rangle|=F(g)$, can be analytically calculated using the conformal perturbation theory [83, 63]. Using the conformal perturbation formalism developed in Ref. [63], the leading correction to the overlap is given by
$|\langle\mathcal{C}_{+}|\Psi_{0}(g)\rangle|\equiv F(g)=2^{1/4}+\frac{\pi\mathcal{A}}{2}\cdot g+\mathcal{O}\left(g^{2}\right)\,,$ (A59)
where $\mathcal{A}=-\sqrt{\frac{2}{3}}$ is the amplitude of the one-point crosscap correlator
$\langle 0,0|\frac{2}{\sqrt{3}}:\mathbf{J}\cdot\bar{\mathbf{J}}:(w,\bar{w})|\mathcal{C}_{+}\rangle=-\sqrt{\frac{2}{3}}\cdot\frac{1}{(1+|w|^{2})^{2}}\equiv\frac{\mathcal{A}}{(1+|w|^{2})^{2}}\,.$ (A60)
Here $|0,0\rangle$ is the vacuum of the $\mathrm{SU}(2)_{1}$ WZW CFT.
The perturbation operator $\frac{2}{\sqrt{3}}:\mathbf{J}\cdot\bar{\mathbf{J}}:$ in Eq. (A57), with conformal weight $h=\bar{h}=1$, is marginal at first order. The one-loop renormalization group (RG) equation, derived by considering an infinitesimal scale transformation $a\to(1+\delta l)a$ ($a$: short distance cutoff at a given scale) shows that the perturbation is indeed marginal irrelevant [84, 70]:
$\frac{\mathrm{d}g}{\mathrm{d}l}=-\pi bg^{2}+\mathcal{O}(g^{3})\,,$ (A61)
where $b=4/\sqrt{3}$ is determined by the three-point correlator on the plane
$\left\langle\frac{2}{\sqrt{3}}:\mathbf{J}\cdot\bar{\mathbf{J}}:(w_{1},\bar{w}_{1})\cdot\frac{2}{\sqrt{3}}:\mathbf{J}\cdot\bar{\mathbf{J}}:(w_{2},\bar{w}_{2})\cdot\frac{2}{\sqrt{3}}:\mathbf{J}\cdot\bar{\mathbf{J}}:(w_{3},\bar{w}_{3})\right\rangle=\frac{-b}{|w_{1}-w_{2}|^{2}|w_{1}-w_{3}|^{2}|w_{2}-w_{3}|^{2}}\,.$ (A62)
To account for the finite-size correction arising from the marginally irrelevant perturbation, we consider a spin-1/2 Heisenberg chain with $N$ sites and lattice spacing $a$ (short-distance cutoff at the ultraviolet scale), with periodic boundary condition imposed. The RG flow is from a theory at the ultraviolet (UV) scale with microscopic coupling $g_{\mathrm{uv}}$ to its low-energy effective description [Eq. (A56)] at the infrared (IR) scale with effective coupling $g$. The length of the system $L=Na=N^{\prime}a^{\prime}$ is fixed under the RG flow, $e^{l}=\frac{a^{\prime}}{a}=\frac{N}{N^{\prime}}$, where $a^{\prime}$ is the short-distance cutoff at the IR scale. Solving the RG equation [Eq. (A61)] gives the effective coupling $g$ as the function of $N$:
$g\equiv g(l)=\frac{g_{\mathrm{uv}}}{1+\pi bg_{\mathrm{uv}}l}=\frac{g_{\mathrm{uv}}}{1+\pi bg_{\mathrm{uv}}\ln(N/N^{\prime})}\,.$ (A63)
Specifically, in the limit $\ln N\gg 1/g_{\mathrm{uv}}$, the asymptotic finite-size scaling behavior of the effective coupling $g$ is independent of microscopic details [84, 70]:
$g\to\frac{1}{\pi b\ln N}\,.$ (A64)
This leads to the universal marginal correction of the crosscap overlap [Eq. (A59)]:
$|\langle\mathcal{C}_{+}|\Psi_{0}(g)\rangle|=F(g)\to F(1/\pi b\ln N)=2^{1/4}-\frac{1}{4\sqrt{2}}\cdot\frac{1}{\ln N}+\mathcal{O}\left(1/(\ln N)^{2}\right)\,.$ (A65)
As discussed above, the universal finite-size correction due to the marginal term is valid for $\ln N\gg 1/g_{\mathrm{uv}}$. However, in lattice simulations, the total number of sites $N$ is typically not sufficiently large to reach that regime. Nevertheless, the form of Eq. (A59) suggests that the lattice crosscap overlap can be expanded in terms of $1/\ln N$, i.e., $|\langle\mathcal{C}_{\mathrm{latt}}|\Psi_{0}(N)\rangle|=a_{1}+a_{2}/\ln N$, where $|\Psi_{0}(N)\rangle$ is the ground state of the Heisenberg chain and $a_{1}$ and $a_{2}$ are fitting parameters. The fitted Luttinger parameter is given by $K=1/a_{1}^{4}$ (theoretical value for the $\mathrm{SU}(2)_{1}$ WZW CFT: $K=1/2$).